Integration by Partial Fractions


1.) ∫dy/(y^2+2y)


2.) ∫(x^3+x+2)/(x^2-1)dx



1
Expert's answer
2021-05-07T12:32:40-0400

(a)

"\int \frac{dy}{y^2+2y}\newline
=\int \frac{dy}{y(2+y)}"

By partial fraction,

"\frac{1}{y(2+y)}=\frac{A}{y}+\frac{B}{2+y}
\newline
\text{by solving, we get A=1/2 and B= -1/2}"

"\int \frac{dy}{y^2+2y}\newline
=\frac{1}{2}\int (\frac{1}{y}-\frac{1}{2+y})dy\newline
=\frac{1}{2}[ logy-log(2+y)]\newline
=\frac{1}{2} log\frac{y}{(2+y)}"


(b)

By using partial fraction and algebraic indentities,

"\frac{x^3+x+2}{x^2-1}=\frac{(x+1)(x^2-x+2)}{(x+1)(x-1)}=\frac{x^2-x+2}{x-1}=x+\frac{2}{x-1} \newline"

Then, the integral

"\int \frac{x^3+x+2}{x^2-1}dx=\int (x+\frac{2}{x-1} )dx\newline
=x^2+2log(x-1)"